# Sensing with near-infrared laser trapped fluorescent nanodiamonds

### v2

@@ Abstract
Biosensing based on optically trapped fluorescent nanodiamonds is an intriguing research direction potentially allowing to resolve biochemical processes inside living cells. Towards this goal, we investigate near infrared (NIR) laser irradiation at 1064 nm on fluorescent nanodiamonds (FNDs) containing nitrogen-vacancy (NV) centers. The NIR wavelength is a popular choice for optical trapping in biological samples because of its low absorption in bio-samples. By conducting comprehensive experiments, we aim to understand how NIR exposure influences the fluorescence and sensing properties of FNDs and to determine the potential implications for the use of FNDs in various sensing applications. The experimental setup involved exposing FNDs to varying intensities of NIR laser light and analyzing the resultant changes in their optical and physical properties. Key measurements included all-optical fluorescence relaxation, optical spectroscopy, and optically detected magnetic resonance (ODMR) spectra. The findings reveal how increased NIR laser power correlates with alterations in ODMR central frequency but also that charge state dynamics under NIR irradiation of NV centers plays a role. We suggest protocols with NIR and green light that mitigate the effect of NIR, and demonstrate that FND biosensing works well with such a protocol.
@@
preprint: APS/123-QED
I Introduction
Figure 1: Graphical abstract illustrating the sensing modalities explored to investigate effects of NIR laser on NV center for sensing. Temperature, abundance of paramagnetic species, and pH, are altered while optical spectroscopy, ODMR spectra, and fluorescence polarization relaxometry is measured from optically trapped nanodiamonds.
@@ I Introduction
Fluorescent nanodiamonds (FNDs) hosting negatively-charged nitrogen-vacancy (NV-) centers combine bright, photostable fluorescence with quantum-coherent electron-spin states, making them uniquely suited for bio-sensing and bio-imaging at the nanoscale [Schirhagl2014, Liu2021]. Because NV spins can be optically polarised and read out at room temperature, changes in their spin-state populations translate directly into optical signals, enabling quantitative sensing of magnetic fields, temperature and local electro-chemical potentials [Doherty2013, Rondin2014]. The ultimate performance of those sensors, however, is governed by the spin- and charge-state photodynamics of the NV centers themselves; understanding how an experimental control parameter perturbs those dynamics is therefore a prerequisite for accurate, high-sensitivity measurements.
@@
The NV center is an electronic spin-1 system with a ground-state zero-field splitting (ZFS) $D\approx 2.87~\mathrm{GHz}$ that separates the $m_{s}=0$ and $m_{s}=\pm 1$ sub-levels. External magnetic fields lift the degeneracy of the $m_{s}=\pm 1$ levels through the Zeeman effect, while lattice strain, temperature and electric fields shift the ZFS and the optical transition energies via the Stark effect [Maze2008, Dolde2011]. In addition, the relative abundance of the neutral (NV0) and negative (NV-) charge states—and thus the fluorescence spectrum and intensity—depends sensitively on the local Fermi level and on photo-ionisation/recombination pathways. These intertwined dependencies let NV centers act as multi-modal nano-probes, provided that the separate interaction between all external controls (laser beams, microwave fields, etc.) and the FNDs are themselves well understood and do not compromise sensor performance.
@@
A particularly powerful external control is optical trapping. Near-infra-red (NIR) optical tweezers can confine—and manoeuvre—sub-100  nm diamonds inside living cells or complex micro-fluidic environments, offering spatially resolved sensing with a resolution set by the nanoparticle size rather than by the optical diffraction limit [Neukirch2013]. Operating the trap at 1064  nm keeps tissue absorption low and mitigates photo-damage [Svoboda1994], yet the same intense NIR field may also heat the nanodiamond, alter NV charge-state kinetics or modify spin-lattice relaxation. Earlier work on bulk diamond or levitated nanodiamonds indeed reported substantial NIR-induced changes in fluorescence and $T_{\rm relax}$[Aslam2013, Ji2016], but a clear picture for FNDs dispersed in aqueous media—arguably the configuration of greatest practical relevance-is missing.
Here, as illustrated by Fig.1, we address that gap. Using freely diffusing and optically trapped FNDs in water and in biologically relevant buffer solutions, we quantify how a 1064  nm focused trapping beam (30–60  mW average power) affects NV optical and spin properties. We employ tree complementary read-outs:
  1. 1.
    photoluminescence (PL) spectroscopy to monitor charge-state redistribution [Geiselmann2013];
  2. 2.
    optically detected magnetic resonance (ODMR) to track shifts in the resonance frequencies with kHz precision [Gruber1997, Sow2020]; and
  3. 3.
    fluorescence relaxometry to probe magnetic-noise spectra—e.g. from free radicals or $\mathrm{Gd}^{3+}$ ions—at the NV transition frequency [Hall2016].
Together, these observables allow us to disentangle three mechanisms that NIR light can induce:
  1. 1.
    Photothermal heating of the FND, shifting the ZFS via the temperature coefficient $dD/dT\approx-74~\mathrm{kHz\,K^{-1}}$[Acosta2010];
  2. 2.
    Power-dependent ionisation/recombination that re-weights NV- and NV0, thereby changing both ODMR spectral properties and PL brightness [Aslam2013, Ji2016];
  3. 3.
    Modified surface sensitivity, whereby the NIR field—through either heating or altered charge state—amplifies or suppresses the NV response to environmental parameters such as pH or paramagnetic species concentration [Fu2010].
@@
We first benchmark these effects in air-dried FNDs, then in trapped FNDs immersed in solutions of controlled pH and temperature, and finally in the presence of a paramagnetic relaxant ($\mathrm{Gd}^{3+}$). By comparing experiments with and without the NIR trap, we establish quantitative limits on NIR-induced artefacts and delineate operating regimes where accurate, spatially resolved biosensing is possible.
II Methods
II.1 Sample Preparation
@@
Fluorescent FNDs with a nominal diameter of 120 nm (Sigma-Aldrich, 798088) were used for most of the experiments reported in this article if not stated otherwise. They possess a significant number of nitrogen-vacancy (NV) centers due to a high concentration of 3ppm. To prepare the FND suspension, 1 µL of the commercial stock solution of FNDs was diluted in 300-700 µL of deionized water depending on the desired concentration and before sonicating for 10 minutes. Next, the FND suspension was deposited on a petri dish (Mattek, part no P35G-1.5-10-C.), thoroughly cleaned to remove any contaminants that could interfere with the experiments and plasma ionised in order to make the surface hydrophilic, ensuring an even distribution of FNDs on the surface. Plasma treatment was applied for 30 seconds in a simple custom-device with a closed chamber in PMMA and under ambient conditions. Plasma was generated with a BD-20AV high-frequency generator (230 V, ETP, IL, USA) with an electrode positioned close to the petri dish surface. The plasma treatment modifies the surface energy of the petri dish by breaking chemical bonds on the surface and introducing polar functional groups, thereby inducing the hydrophilic state [Al-Azzam2022]. Approximately 3 µL of the FND suspension was then carefully deposited onto the clean petri dish using a micropipette. The suspension was gently spread across the surface before allowing it to dry. The drying process took place at room temperature to minimize any potential alterations to the FND properties potentially caused by excessive heating or cooling. After the solution had completely evaporated, the sample was ready for the optical experiments described in the following. The preparation of FNDs for experiments in the liquid environment also starts with the dilution steps and cleaning of the petri dish, but to maintain the FNDs in solution, the petri dish was not plasma treated and the solution was not left for evaporation.
@@
FNDs with a nominal size of 70 nm were used for sensing experiments as they are preferably used in our most recent cell experiments [Niora2024]. The FNDs (Sigma-Aldrich 1003626896) were prepared following the same protocol. A suspension of FNDs was created by diluting 1 µL of the commercial stock solution in 1000 µL of deionized water, followed by sonication for 10 minutes to ensure a homogeneous dispersion. Approximately 50 µL of the suspension was deposited in a sterile Petri dish that was freshly unwrapped.
II.2 Optical Setup
@@
The dried FND sample was placed inside a mini-incubator (pecon XS 2000) in an optical setup based on a customized Thorlabs OTKB/M modular optical tweezer system. Customization included a different objective and long-distance condenser lens. Vendors and other details are listed in Table 2 in the Appendix. The setup is constructed around an inverted microscope configuration, equipped with a high-resolution camera that capture wide-field bright-field images of the FND, and two lasers for the simultaneous illumination and manipulation of the FNDs. A green laser with a wavelength of 532 nm was utilized for the excitation of the NV centers, enabling the initialization and readout of their spin states. The second laser is operating in the near-infrared (NIR) at a wavelength of 1064 nm and used for FND trapping. This wavelength falls within the biological transparency window, which minimizes absorption and reduces the risk of significant heating effects in biological samples. The two lasers were aligned and their beams combined using a dichroic mirror, ensuring that they shared the same optical path before being focused onto the sample through the microscope objective, as detailed in Fig. 2a). By trapping an FND and then finetuning the position of the green laser to maximize the signal on the photon counter, we ensured the same focal point for both the green and NIR lasers. Additionally, acousto-optic deflectors were used to precisely control the timing and intensity of the laser beams by using Qudi software, enabling fine-tuned manipulation of the FNDs.
@@
As illustrated in Fig. 2a), several modes of detection were installed: i) camera detection, ii) spectrometer detection and iii) photon count detection. A motorized flip-mirror ensures the ability to rapidly shift between camera detection and either spectrometer or photon count detection. The fluorescence spectrum of the FNDs is recorded and analyzed by a fiber coupled spectrometer while alternatively photon counts are measured with a fiber coupled avalanche photo detector. In addition, to measure the CW ODMR spectrum of FNDs, a signal generator connected to an amplifier is connected to a single-turn copper loop antenna (5 mm) placed approximately 1 mm above the sample. When the flip-mirror is positioned towards the camera, the latter captures high-resolution images of the FNDs illuminated during the experiments in real time. The combination of different detection methods thus made it possible to continuously monitor the position of the FNDs, determine whether they were isolated or in aggregates, and analyse their photoluminescence properties throughout the experimental process.
Figure 2: a) Illustration of the optical setup with the most important lasers, optical lenses, filters and equipment. Details of the components are specified in Table 2 in the appendix. b) Laser sequence for fluorescence relaxometry measurement. In green the 532 nm laser and in red the 1064 nm laser. Pulse-widths are $t_{\rm trap}=20\mu s$, $t_{ini}=35\mu s$, $t_{RO}=15\mu s$ while the variable waiting time $\tau$ allows to record the relaxometry at times spanning from $7.2\mu s$ to $720\mu s$.
II.3 FND Trapping
@@
All trapping experiments were performed in a fluid chamber containing 50 µL of the FND aqueous suspension unless stated otherwise. The NIR laser was utilized for trapping the FNDs in the focused beam [Neuman2004, Gennerich2017, Jones2015]. For relaxometry, the temporal sequence of the green and trapping lasers, as illustrated in Fig. 2b), avoids temporal overlap between the two lasers while maintaining the FND stably trapped [Russell2018a]. Strictly speaking, this protocol does not constitute a conventional spin-relaxometry measurement. With an all-optical protocol, the method inherently involves potential inaccuracies due to photoionization induced by the green laser itself, thus affecting the NV- charge states. To address this limitation, measurement protocols incorporating microwave pulses explicitly designed to disentangle spin dynamics from photoionization effects have been developed. Examples include the microwave $\pi$-pulse method introduced by Ref. [Jarmola2023], as well as alternative sequences described in Ref. [Cardoso2023]. We hence will refer to the measured decay time as $T_{\rm relax}$ throughout the remainder of this paper. The green laser was pulsed with initialization time $t_{ini}$ of 35 µs, and read-out time $t_{RO}$ of 15 µs, for the readout of the NV center fluorescence, while the NIR laser was on to trap the FNDs in sequences with duration $t_{\rm trap}$ of 20 µs.
II.4 Sensing Experiments
@@
For the experiments detecting paramagnetic species, a gadolinium solution (0.45 M in acetate) was gradually added to the FND’s environment. The concentrations were incrementally adjusted, and measurements were performed after each addition to evaluate the changes in the NV center properties.
@@
For the pH sensing experiments, an initial 0.1 M NaOH(aq) solution at pH 13 was prepared. The pH was then progressively decreased by adding controlled volumes of a 1 M HCl solution. After each pH adjustment, the properties of the NV centers were measured while the FNDs were maintained in the optical trap at 60 mW.
@@
For the sensing of paramagnetic species and pH, changes to the surrounding medium was involved. To ensure that the FND remained trapped despite disturbances in the medium, the CW NIR laser power was temporarily increased to 100 mW while liquid was added. Subsequently, the laser power was reduced back to 60 mW for measurements with the new conditions, and the same FND.
III Results
We begin by presenting the main results of our characterisation experiments, starting with the examination of the general fluorescence properties of NV-centers within the FNDs through photoluminescence (PL) spectroscopy. This foundational assessment provides an overview of the baseline optical characteristics and helps identify any immediate changes induced by the NIR laser irradiation. Following the PL spectrum analysis, we present the variations in fluorescence intensity as a function of optically detected magnetic resonance (ODMR) conditions. This step is crucial for understanding the impact of NIR irradiation on the magnetic resonance properties of the NV centers. Finally, we conduct fluorescence relaxometry to investigate the spin relaxation dynamics under different NIR laser parameters. PL spectrum measurement and ODMR measurement were conducted with CW NIR laser irradiation.
The experiments conducted under dry conditions primarily serve as a thermal control benchmark. This dry setup facilitates the understanding of thermal phenomena, such as localized heating effects, by providing a clear reference to compare with conditions involving FNDs suspended in aqueous environments. We acknowledge that the dry condition itself lacks direct practical application in sensing contexts.
For each set of experiments described in the subsections below, we exposed single FNDs to CW NIR laser at power levels of 0 mW, 30 mW, and 60 mW.
III.1 Photoluminescence Spectrum Analysis
As the NV neutral and negative charge states have distinct optical emission spectra, we conducted PL spectroscopy of FNDs at varying NIR power to investigate its impact. The FNDs were either dried on glass substrates or trapped in water to examine the influence of the surrounding medium. We record the PL spectrum in the range between 550 nm to 745 nm with the spectrometer for each of the different values of the CW NIR laser power.
Figure 3: Effect of CW NIR laser on dry FND emission spectrum. Optical emission spectrum of one FND, dry on glass, under CW NIR laser exposure at 0 mW (green), 30 mW (orange) and 60 mW (red) of 1064 nm laser power.
Upon CW NIR laser light exposure, our PL measurements revealed a significant drop in the total PL intensity. Notably, as a function of NIR laser power there was a slightly more pronounced reduction in the spectral intensity of NV- compared to NV0. For the fluorescence spectrum (like Fig. 3), the curves remain quite similar when water is added, although with less effect of increase in NIR laser intensity from 30mW to 60mW (data not shown) and the difference between 30 mW and 60 mW is even more reduced for FNDs that are trapped in water.
To quantify the changes in relative charge state emission intensities under CW NIR laser irradiation, we determine the ratio $\sigma_{0-}$ between NV0 and NV- emission intensities by integrating over the spectral ranges for NV0 (550–620 nm) and NV- (655–745 nm), respectively. Compared to the analysis used in Ref. [Sow2020], we applied a normalization factor 19/14 to account for the difference in the relative spectral ranges:
For FNDs trapped with NIR powers of 30 mW and 60 mW, we observe an averaged similar emission ratios $\sigma_{0-}$ of $0.48\pm 0.09$. In contrast, for FNDs dried on glass, we measure an emission ratio of $0.43\pm 0.11$ without NIR irradiation that increases to $0.57\pm 0.15$ for a NIR laser power of 60 mW NIR. All results are displayed in Fig. 6 parts a) and c).
III.2 Optically Detected Magnetic Resonance Measurements
In the next step, we elaborate on the influence of CW NIR laser light on the ODMR resonance frequency and contrast, revealing the impact of the NIR trapping laser on the NV electron spin properties.
The ODMR spectra were acquired in continuous-wave (CW) mode with constant intensities of both the 532 nm pump and 1064 nm trapping lasers and the microwave source, while sweeping the frequency of the microwave in steps of 0.3 MHz. Generally and for all FNDs, we observe a redshift in the resonance frequency with increasing CW NIR laser power, as illustrated in Fig. 4 by the spectra from one FND with 0 mW and 60 mW NIR power.
Figure 4: Effect of CW NIR laser power on ODMR spectrum of dry FND. ODMR spectra of one FND under CW NIR laser exposure at 0 mW (green) and 60 mW (red) of 1064 nm laser power. The resonance frequency shift and decrease in contrast with increasing CW NIR power is observed. The power of the 532 nm laser was 0.7 mW, measured before the objective lens, and the microwave amplifier was set to -2dB.
For FNDs dried on glass substrates, the resonance frequency shifted significantly to the lower frequencies. Taking 2.87 GHz as reference value for the zero field splitting, the resonance frequency shift increased in absolute value from $0.31\pm 0.54$ MHz at 0 mW to $5.83\pm 046$ MHz at 60 mW of CW NIR power. In contrast, FNDs trapped in water exhibited much smaller frequency shifts, with changes of $0.84\pm 0.21$ MHz at 30 mW and $1.58\pm 0.33$ MHz at 60 mW. Examples are shown in Fig. 4 and combined results displayed in Fig. 7. As before, for the similar experiment when water was added to the same FND that remained stuck on the glass slide, the ODMR spectra show a trend very similar to Fig. 4 although the difference between 30 mW and 60 mW of NIR power is less pronounced (data not shown).
[…]
III.3 Relaxometry Measurements
Last, we conducted relaxometry experiments under different NIR laser power, to extract values for the corresponding $T_{\rm relax}$. As above, FNDs were either dried on glass substrates or suspended in water to assess the influence of the surrounding medium on the relaxometry measurements with NIR irradiation. For these experiments, the all-optical pulsing sequence in Fig. 2b) was applied.
[…]
Figure 5: Effect of NIR laser power on $T_{\rm relax}$ relaxation times of dry FNDs. $T_{\rm relax}$ relaxometry measurements for one FND dried on glass under NIR laser powers of 0 mW (green) and 60 mW (red) at 1064 nm wavelength. Pulsing as in Fig 2b).
[…]
Overall, as showed in Fig. 6 and Fig. 7, the data suggest that the aqueous environment reduces the effect of NIR irradiation on ODMR-frequency shifts. Further, our results confirm that with specific NIR irradiation sequences and optimizing laser parameters, cf. Fig. 2b) and Ref. [Russell2018a], potential detrimental effects of NIR on $T_{\rm relax}$ relaxation can be significantly mitigated.
Figure 6: a) Correlation between NIR laser power, ODMR resonance frequency shift and $\sigma_{0-}$ fluorescence intensity ratio for FNDs trapped inside water. b) Correlation between NIR laser power, ODMR resonance frequency shift and fluorescence relaxometry times for FNDs trapped inside water. c) Correlation between NIR laser power, ODMR resonance frequency shift and $\sigma_{0-}$ fluorescence intensity ratio for FNDs dry on glass. d) Correlation between NIR laser power, ODMR resonance frequency shift and fluorescence relaxometry times for FNDs dry on glass. Data for 30 FNDs exposed to 0 mW (green), 30 mW (orange), and 60 mW (red) of 1064 nm NIR laser irradiation. CW NIR for ODMR and spectra, pulsed NIR as in Fig 2b) for fluorescence relaxometry. In general, we observe what appears to be a trend with differences between the three powers of NIR laser light for FNDs dry on glass, and and no or much less pronounced trend for FNDs in the optical trap in water.
Figure 7: Summary of experimental results, showing variation for individual nanodiamonds. (a) ODMR resonance frequency shift, relative to the expected resonance at 2.87 GHz, (b) fluorescence intensity ratio $\sigma_{0-}$, and (c) $T_{\rm relax}$ relaxometry times as functions of NIR laser power. Data for 5 random FNDs dry on glass (black stars) and trapped in water (gray circles) with 1064 nm NIR laser irradiation.
Table 1: ANOVA test results for the effect of NIR laser power on FNDs dried on glass substrates, on glass after adding water and trapped in water. The p-values indicate the significance of differences observed across NIR power levels for $T_{\rm relax}$ relaxation times, ODMR frequency shifts, and $\sigma_{0-}$ ratios.
p-value
FNDs dried on glass
$T_{\rm relax}$ Relaxation Time <0.0001<0.0001
ODMR Frequency Shift <0.0001<0.0001
$\sigma_{0-}$ Ratio 0.0024
FNDs on glass with water
$T_{\rm relax}$ Relaxation Time $0.072$
ODMR Frequency Shift $0.008$
$\sigma_{0-}$ Ratio 0.201
FNDs Trapped in Water
$T_{\rm relax}$ Relaxation Time 0.221
ODMR Frequency Shift 0.073
$\sigma_{0-}$ Ratio 0.371
For FNDs dried on glass substrates, the very low p-values (p <0.01<0.01) for all three parameters confirm that the differences observed with varying NIR laser powers are statistically significant. This indicates that NIR irradiation has a considerable impact on the NV centers’ properties in air.
For FNDs on glass substrates after adding water, the mean p-values ($0.01<p<0.2$) for all three parameters confirm that the differences observed with varying NIR laser powers are not always statistically significant. In fact, the difference between no NIR laser and NIR laser is significant but the differences between 30 mW and 60 mW power of NIR laser are not significant.
In contrast, for FNDs trapped in water, the p-values are high (p >0.07>0.07), suggesting that the differences observed across different NIR power levels are not statistically significant. This reinforces the conclusion that the aqueous environment mitigates the effects of NIR irradiation on the NV centers.
III.4 Model of Charge State Dynamics

### v3

@@ Abstract
Biosensing based on optically trapped fluorescent nanodiamonds potentially allows to resolve biochemical processes inside living cells at a desired intracellular location. Towards this goal, we investigate near infrared (NIR) laser irradiation at 1064 nm on fluorescent nanodiamonds (FNDs) containing nitrogen-vacancy (NV) centers. The 1064 nm NIR wavelength is a popular choice for optical trapping because of its low absorption in bio-samples. By conducting comprehensive experiments, we aim to understand if and how NIR exposure influences the fluorescence and sensing capabilities of FNDs and to determine the potential implications for the use of FNDs in various sensing applications. Our experiments exposed FNDs to varying intensities of NIR laser light while carefully monitoring their optical and magnetic properties. Key measurements included all-optical fluorescence relaxation, optical spectroscopy, and optically detected magnetic resonance (ODMR) spectra. The findings reveal how increased NIR laser power correlates with alterations in ODMR central frequency but also that charge state dynamics under NIR irradiation of NV centers play a role. We demonstrate that FND biosensing works well with a protocol involving both NIR and green light, while mitigating the effect of NIR.
keywords:
@@
journal: Diamond and related materials
[…]
{graphicalabstract}[Uncaptioned image]
{highlights}
Demonstration of multi-modal NV-center based sensing with optically trapped nanodiamonds
Exploring the effect of a 1064 nm NIR trapping laser on the NV-center readouts
1 Introduction
@@
Fluorescent nanodiamonds (FNDs) hosting negatively-charged nitrogen-vacancy (NV-) centers combine bright, photostable fluorescence with quantum-coherent electron-spin states, making them uniquely suited for bio-sensing and bio-imaging at the nanoscale Schirhagl et al. (2014); Hsiao et al. (2016). Because NV spins can be optically polarised and read out at room temperature, changes in their spin-state populations translate directly into optical signals, enabling quantitative sensing of magnetic fields, temperature and local electro-chemical potentials Doherty et al. (2013); Rondin et al. (2014). The ultimate performance of those sensors, however, is governed by the spin- and charge-state photodynamics of the NV centers themselves; understanding how an experimental control parameter perturbs those dynamics is therefore a prerequisite for accurate, high-sensitivity measurements.
@@
The negatively charged NV center is an electronic spin-1 system with a ground-state zero-field splitting (ZFS) $D\approx 2.87~\mathrm{GHz}$ that separates the $m_{s}=0$ and $m_{s}=\pm 1$ sub-levels. External magnetic fields lift the degeneracy of the $m_{s}=\pm 1$ levels through the Zeeman effect, while lattice strain, temperature and electric fields shift the ZFS and the optical transition energies via the Stark effect Maze et al. (2008); Dolde et al. (2011). In addition, the relative abundance of the neutral (NV0) and negative (NV-) charge states—and thus the fluorescence spectrum and intensity—depends sensitively on the local Fermi level and on photo-ionisation/recombination pathways. These intertwined dependencies let NV centers act as multi-modal nano-probes of their local environment, provided that the separate interaction between all external controls (laser beams, microwave fields, etc.) and the FNDs are themselves well understood and do not compromise sensor performance.
@@
A particularly powerful external control is optical trapping. Near-infra-red (NIR) optical tweezers can confine and manoeuvre sub-100  nm diamonds inside living cells or complex micro-fluidic environments, offering spatially resolved sensing with a resolution set by the nanoparticle size rather than by the optical diffraction limit Horowitz et al. (2012). Operating the trap at 1064  nm keeps tissue absorption low and mitigates photo-damage Svoboda and Block (1994), yet the same intense NIR field may also heat the nanodiamond, alter NV charge-state kinetics or modify spin-lattice relaxation Lai et al. (2013); Geiselmann et al. (2013); Ji and Dutt (2016). Earlier work on bulk diamond or levitated nanodiamonds indeed reported substantial NIR-induced changes in fluorescence and the spin relaxation $T_{\rm relax}$Neukirch et al. (2013); Ji and Dutt (2016), but a clear picture for FNDs dispersed in aqueous media—arguably the configuration of greatest practical relevance-is missing.
Here, as illustrated by the graphical abstract, we systematically address that gap. Using freely diffusing and optically trapped FNDs in water and in biologically relevant buffer solutions, we quantify how a 1064 nm focused trapping beam with an average power between 30–60 mW focused to a spot with an Airy diameter less than 450 nm affects NV optical and spin properties. We employ three complementary read-outs, each indicative of a particular NV property: i) photoluminescence (PL) spectroscopy to monitor charge-state redistribution between NV0 and NV-Aslam et al. (2013); Geiselmann et al. (2013); Sow et al. (2020); ii) optically detected magnetic resonance (ODMR) at zero bias field to track shifts in the resonance frequencies with kHz precision which we discuss as a measure of temperature change  Gruber et al. (1997); Acosta et al. (2010); Doherty et al. (2014); Hall et al. (2016); and iii) fluorescence relaxometry to probe magnetic-noise spectra—e.g. from free radicals or $\mathrm{Gd}^{3+}$ ions—at the NV transition frequency Tetienne et al. (2013); Mzyk et al. (2022). Together, these observables allow us to address and potentially disentangle three mechanisms that NIR light has been suggested to induce, namely a) Photothermal heating of the FND, shifting the ZFS via the temperature coefficient $dD/dT\approx-74~\mathrm{kHz\,K^{-1}}$Acosta et al. (2010); Lai et al. (2013); b) Power-dependent ionisation/recombination that re-weights NV- and NV0, thereby changing both ODMR spectral properties and PL brightness Aslam et al. (2013); Ji and Dutt (2016); and c) Modified surface sensitivity, whereby the NIR field—through either heating or altered charge state—amplifies or suppresses the NV response to environmental parameters such as pH or paramagnetic species concentration Fu et al. (2010).
@@
Initially, we compare these effects in air-dried FNDs, then in trapped FNDs immersed in an aqueous solution. Then, to address potential impact of NIR on sensing capabilities of the FNDs, we measure on trapped FNDs in solutions of controlled pH and temperature, and finally in the presence of a paramagnetic relaxant ($\mathrm{Gd}^{3+}$). Thereby, we establish quantitative limits on NIR-induced artefacts and delineate operating regimes where accurate, spatially resolved biosensing is possible.
2 Methods
2.1 Sample Preparation
@@
Fluorescent FNDs with a nominal diameter of 120 nm (Sigma-Aldrich, 798088) and an NV concentration of 3ppm were used for most of the experiments reported in this article if not stated otherwise. To prepare the FND suspension, 1 µL of the commercial stock solution was diluted in 300-700 µL of deionized water depending on the desired concentration and before sonicating for 10 minutes. Next, the petri dish (Mattek, part no P35G-1.5-10-C) was prepared before deposition of FNDs. For this, the petri dish was thoroughly cleaned and plasma ionised in order to make the surface hydrophilic, ensuring an even distribution of FNDs on the surface. Plasma treatment was applied for 30 seconds in a simple custom-device with a closed chamber in PMMA and under ambient conditions. Plasma was generated with a BD-20AV high-frequency generator (230 V, ETP, IL, USA) with an electrode positioned close to the petri dish surface. Approximately 3 µL of the FND suspension was then carefully deposited onto the clean petri dish using a micropipette and the sample was left to dry at room temperature. The preparation of FNDs for experiments in the liquid environment also starts with the dilution steps and cleaning of the petri dish, but to maintain the FNDs in solution, the petri dish was not plasma treated and the solution was not left for evaporation.
@@
FNDs with a nominal size of 70 nm were used for sensing experiments as they are preferably used in our most recent cell experiments Niora et al. (2024); Mzyk et al. (2025, 2026). The 70 nm FNDs (Sigma-Aldrich 1003626896) were prepared following the same protocol. A suspension of FNDs was created by diluting 1 µL of the commercial stock solution in 1000 µL of deionized water, followed by sonication for 10 minutes to ensure a homogeneous dispersion. Approximately 50 µL of the suspension was deposited in a sterile Petri dish that was freshly unwrapped.
2.2 Optical Setup
@@
The dried FND sample was placed inside a mini-incubator (pecon XS 2000) in an optical setup based on a customized Thorlabs OTKB/M modular optical tweezer system. Customization included a different trapping laser, and different objective and long-distance condenser lens. Vendors and other details are provided in the Appendix, Fig. 9 and Table 2. The setup is constructed around an inverted microscope configuration, equipped with white light illumination and a high-resolution camera that capture wide-field bright-field images of the sample, and two lasers for the simultaneous illumination and manipulation of the FNDs. A 532 nm laser was utilized for the excitation of the NV centers, while a 1064 nm near-infrared (NIR) laser was used for FND trapping. The two lasers were aligned and their beams combined using a dichroic mirror and collectively focused onto the sample through the microscope objective, as detailed in Fig. 1, i.e., a single objective setup. By trapping an FND and then finetuning the position of the green laser to maximize the signal on the photon counter, we ensured the same focal point for both the green and NIR lasers. Additionally, acousto-optic deflectors, controlled via a Qudi Binder et al. (2017) interface, were used to precisely control the timing and intensity of the laser beams.
@@
As illustrated in Fig. 1, several modes of detection were installed: i) camera detection, ii) spectrometer detection and iii) photon count detection. A motorized flip-mirror allows to rapidly shift between camera detection and either spectrometer or photon count detection, for recording of FND fluorescence. In addition, to measure the CW ODMR spectrum of FNDs, a signal generator together with an amplifier are connected to a single-turn copper loop antenna (5 mm) placed approximately 1 mm above the sample.
Figure 1: Essential elements of the optical setup. NV centers are excited with 532 nm green light, and nanodiamonds are optically trapped with the 1064 nm NIR light. Detection modalities include bright-field, wide-field imaging on a camera as well as photon counting and fluorescence spectroscopy. Zoom on the sample shows a) the situation with NIR and green light focused on an FND on glass, in dry conditions. Similarly b) shows the same but with added water while c) illustrates the FND being optically trapped in aqueous medium. d) Laser sequence for fluorescence relaxometry measurement. In green the 532 nm laser and in red the 1064 nm laser. Pulse-widths are $t_{\rm trap}=20\mu s$, $t_{ini}=35\mu s$, $t_{RO}=15\mu s$ while the variable waiting time $\tau$ allows to record the relaxometry at times spanning from $7.2\mu s$ to $720\mu s$.
2.3 FND Trapping
@@
All trapping experiments were performed in a fluid chamber containing 50 µL of the FND aqueous suspension unless stated otherwise. The NIR laser was utilized for trapping the FNDs in the focused beam Neuman and Block (2004); Jones et al. (2015); Gennerich (2017). FND trapping was confirmed through standard analysis of histograms and power spectral density of the positions of the FND Berg-Sørensen and Flyvbjerg (2004), with positions recorded by back focal plane interferometry on a quadrant photodiode Allersma et al. (1998). Examples are shown in the Supplementary information. For relaxometry, the temporal sequence of the green and trapping lasers, as illustrated in Fig. 1d), avoids temporal overlap between the two lasers while maintaining the FND in the trap Russell et al. (2018). Strictly speaking, this protocol does not constitute a conventional spin-relaxometry measurement. We hence will refer to the measured decay time as $T_{\rm relax}$ throughout the remainder of this paper. The green laser was pulsed with NV initialization time $t_{ini}$ of 35 µs and read-out time $t_{RO}$ of 15 µs, while the NIR laser was turned on with duration $t_{\rm trap}$ of 20 µs to trap the FNDs.
2.4 Sensing Experiments
@@
For the experiments detecting paramagnetic species, a gadolinium solution (0.45 M in acetate) was gradually added to the FND’s environment. The concentrations were incrementally adjusted, and measurements on FNDs were performed after each addition to evaluate the changes in the NV center properties.
@@
For the pH sensing experiments, an initial 0.1 M NaOH(aq) solution at pH 13 was prepared. The pH was then progressively decreased by adding controlled volumes of a 1 M HCl solution. After each pH adjustment, the properties of the NV centers were measured while the FNDs were maintained in the optical trap with a power of 60 mW, corresponding to an intensity of order $10^{7}$ W/cm2.
@@
For the sensing of paramagnetic species and pH, changes to the surrounding medium was involved. To ensure that the FND remained trapped despite disturbances in the medium, the CW NIR laser power was temporarily increased to around 100 mW while liquid was added. Subsequently, the laser power was reduced back to 60 mW for measurements with the new conditions, and the same FND.
2.5 Simulations - Model of Charge State Dynamics